OpenStudy (anonymous):

Find the slope-intercept equation of a line given the conditions. The slope is 5 and the y-intercept is (0,9). Please explain the steps you took to get your answer.

OpenStudy (anonymous):

y = 5x+9

OpenStudy (saifoo.khan):

y = 5x +9

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

you use the point slope equation y-y1=m(x-x1) y-9=5(x-0) y=5x+9

OpenStudy (anonymous):

Thank you guys

OpenStudy (anonymous):

no problem